OpenStudy (anonymous):

graph the system of linear inequalities. Then state two possible solutions. y> 2x + 2 y< -3/2x - 5

OpenStudy (raden):

can u graph a straight line equation y=mx+c ?

OpenStudy (anonymous):

No we just started this unit.

OpenStudy (anonymous):

first y=2x+2 y=-3/2x-5

OpenStudy (anonymous):

Okay, I'm following along so far.

OpenStudy (anonymous):



OpenStudy (anonymous):

and u make it anotther equation like me ok.

OpenStudy (richyw):

so anything in the dark red would satisfy both inequalities

OpenStudy (richyw):

the blue is \(y>2x+2\) and the red is \(y<−\frac32 x−5\)

OpenStudy (richyw):

so all you have to do is graph the lines \(y=2x+2\) and \(y=-\frac32 x-5\) (I would use a dotted line for \(>\) or \(<\) and a solid line for \(\geq\) or \(\leq\)). Then shade in above or below the line depending what is required. Then your solutions will look like that graph, and where the shading overlaps is where you get a solution